HETAV PATEL

Grade 11,

the specific heat of a substance varies with temperature according to c=0.20+0.14T+0.023T²,with T in °C and c in cal÷kg K find the energy required to raise the temperature of 2.0g of this substance from 5.0°C to 15°C.

the specific heat of a substance varies with temperature according to c=0.20+0.14T+0.023T²,with T in °C and c in cal÷kg K find the energy required to raise the temperature of 2.0g of this substance from 5.0°C to 15°C.

Grade:12

1 Answers

Shaik Irfan Kareem Ali
10 years ago
q=mc(t2-t1) c=integrate c(T) from 5 to 15 c=40.92cal/kg q=0.002(40.92)(10) q=0.82 cal